OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

OpenStudy (anonymous):

@nikato do you think you can help me again? :)

OpenStudy (nikato):

So your equation x-62599=0.25y Substitute 125000 for y because that's the # of people who voted. And now you should be able to solve for x

OpenStudy (anonymous):

O crap I have no clue :( ummmm....

OpenStudy (nikato):

So you'll have x-62500=0.25(125000)

OpenStudy (anonymous):

that's the answer or the equation to solve it? I'm sorry I'm horrible with algebra

OpenStudy (calculusxy):

@LilliBelle

OpenStudy (calculusxy):

You need to solve for x in the equation that nikato provided.

OpenStudy (anonymous):

Okay that's what I thought I just don't know how to go about it haha I've taken algebra 5 times now I know that makes me sound stupid but its just im rlly rlly rlly bad at algebra

OpenStudy (calculusxy):

No problem. You'll get the hang of it :)

OpenStudy (calculusxy):

Do now understand on how to solve it?

OpenStudy (anonymous):

....no :(

OpenStudy (calculusxy):

I will ask you a few questions and you will answer it with as much as you know :)

OpenStudy (calculusxy):

How many total votes were casted?

OpenStudy (anonymous):

125000?

OpenStudy (calculusxy):

Yes

OpenStudy (calculusxy):

How much (percentage) did Maya receive?

OpenStudy (anonymous):

ummm 75%?

OpenStudy (calculusxy):

No 25% if you look carefully.

OpenStudy (calculusxy):

Using these two pieces of info, we can then try to find .25 of 125,000 then add that product to 62500.

OpenStudy (anonymous):

okay do divide 25% from 125,000 and then when I get the answer add it to 62500

OpenStudy (calculusxy):

No multiply

OpenStudy (calculusxy):

.25 * 125,000

OpenStudy (calculusxy):

Then add.

OpenStudy (anonymous):

So the answer is 93,750